# High Dimensional Ensemble Kalman Filter

### v1

@@ 6.1 Lorenz-96 model
Panels (d), (g)-(f) of Figure 2 demonstrate the influence of the ensemble size on the analysis RMSEs, which show that the analysis RMSEs of the inflation and the high dimensional EnKF methods became smaller with the increase of the ensemble size $n$. The proposed high dimensional EnKF methods performed much better than the inflation method with smaller RMSEs, especially when the ensemble size $n$ was small (panels (d) and (g) with $n=20$). This suggests that the proposed high dimensional EnKF methods could provide more accurate analysis states when the ensemble size was not big enough. Ensemble size is a significant restriction for the running of large geophysical models as they are extremely costly to run.
Table 1: Average RMSEs to the oracle analysis state and true state of the five EnKF schemes for Lorenz-96 model with $p=q=40,100$, $n=30$, $\sigma_{0}=0$ and $F=8$ for the correct and other $F$ values for the misspecified models. The numbers inside the parentheses are the $25\%$ and $75\%$ quantiles of the RMSEs.
p type method F=6 F=7 F=8 F=9 F=10
$40$ oracle standard 4.27 (4.22,4.33) 4.37 (4.30,4.44) 4.62 (4.54,4.69) 4.92 (4.85,4.99) 5.23 (5.17,5.30)
inflation 1.08 (1.06,1.10) 0.83 (0.81,0.84) 0.60 (0.58,0.62) 0.93 (0.91,0.94) 1.21 (1.18,1.23)
banding 0.88 (0.87,0.88) 0.75 (0.74,0.76) 0.69 (0.68,0.70) 0.79 (0.78,0.80) 0.96 (0.95,0.97)
tapering 0.88 (0.87,0.88) 0.75 (0.74,0.75) 0.68 (0.67,0.69) 0.80 (0.79,0.81) 0.97 (0.96,0.99)
threshold 0.91 (0.90,0.92) 0.78 (0.77,0.79) 0.51 (0.38,0.57) 0.82 (0.81,0.83) 0.98 (0.97,0.99)
true standard 4.31 (4.25,4.36) 4.39 (4.32,4.47) 4.62 (4.54,4.69) 4.91 (4.84,4.98) 5.22 (5.16,5.29)
inflation 1.18 (1.16,1.20) 0.88 (0.86,0.89) 0.59 (0.57,0.62) 0.96 (0.94,0.98) 1.27 (1.25,1.29)
banding 0.98 (0.97,0.99) 0.80 (0.79,0.81) 0.71 (0.70,0.72) 0.83 (0.82,0.84) 1.04 (1.03,1.06)
tapering 0.98 (0.97,0.99) 0.80 (0.79,0.81) 0.70 (0.69,0.71) 0.84 (0.83,0.85) 1.06 (1.05,1.07)
threshold 1.02 (1.01,1.03) 0.83 (0.82,0.84) 0.57 (0.45,0.62) 0.87 (0.86,0.88) 1.07 (1.06,1.08)
$100$ oracle standard 4.40 (4.37,4.44) 4.58 (4.54,4.62) 4.84 (4.80,4.88) 5.14 (5.10,5.18) 5.45 (5.41,5.49)
inflation 2.91 (2.88,2.95) 2.83 (2.79,2.88) 2.88 (2.83,2.92) 3.11 (3.05,3.15) 3.40 (3.36,3.44)
banding 1.15 (1.13,1.16) 0.74 (0.73,0.75) 0.50 (0.49,0.50) 0.89 (0.87,0.91) 1.68 (1.63,1.72)
tapering 1.27 (1.24,1.30) 0.75 (0.74,0.76) 0.46 (0.45,0.47) 1.08 (1.05,1.11) 2.56 (2.46,2.65)
threshold 1.72 (1.65,1.79) 1.04 (1.02,1.06) 0.83 (0.82,0.85) 1.15 (1.12,1.18) 1.89 (1.82,1.96)
true standard 4.43 (4.40,4.46) 4.59 (4.55,4.63) 4.84 (4.80,4.88) 5.13 (5.10,5.17) 5.44 (5.40,5.47)
inflation 2.94 (2.91,2.98) 2.85 (2.80,2.89) 2.87 (2.83,2.92) 3.10 (3.05,3.15) 3.41 (3.37,3.45)
banding 1.26 (1.25,1.28) 0.84 (0.83,0.85) 0.60 (0.59,0.60) 0.97 (0.95,0.99) 1.75 (1.71,1.79)
tapering 1.38 (1.35,1.41) 0.85 (0.84,0.86) 0.57 (0.56,0.58) 1.15 (1.12,1.18) 2.60 (2.51,2.69)
threshold 1.76 (1.70,1.84) 1.06 (1.04,1.08) 0.82 (0.80,0.84) 1.15 (1.12,1.19) 1.92 (1.85,1.99)
Table 1 reports the average analysis RMSEs to the oracle analysis state and true state and their $25\%$ and $75\%$ quantiles by the standard EnKF, the inflation and the high dimensional EnKF methods using the banding and tapering and the thresholding estimators when $\sigma_{0}=0$ which is the same setting as Liang et al. (2012). One can see that the RMSEs to the oracle analysis state and true state of the proposed high dimensional methods were much smaller than the standard EnKF and the inflation method under both the correct and misspecified models. In the case of a higher dimensional setting with $p=100$, the effectiveness of the proposed high-dimensional EnKF schemes became even more pronounced. In contrast, the inflation method demonstrated significantly higher analysis RMSEs, even when applied to the correct model ($F=8$). For instance, when $p=100,F=8$, the banding method achieved an RMSE of 0.5, while the inflation method resulted in a much higher RMSE of 2.88. Besides, the high dimensional EnKF with banding estimator performed the best with the smallest RMSEs in almost all considered settings. Consistent with the theoretical results, the RMSEs to the true state were largely higher than those to the oracle analysis state.
@@
Table 2: Average RMSEs to the oracle analysis state and true state of the five EnKF schemes and their divergence rates (in the parentheses) for Lorenz-96 model for $p=40$, $q=30$, $n=30$, $\sigma_{0}=0.1$, and $F=8$ for the correct and other $F$ values for the misspecified models.
@@
Table 2 displays the analysis RMSEs of different assimilation methods and their divergence rates (in the parentheses) with sparser observations $q=30$ while $p=40$, which illustrates the influence of the sparsity of the observations on the assimilation accuracy. We see that the analysis RMSEs of different assimilation methods were slightly higher than those under more observations $p=q=40$. The RMSEs of the high dimensional EnKF methods were smaller than those of the inflation method and the standard for imperfect models ($F\neq 8$) and were comparable to the inflation method for the correct model ($F=8$). However, it is noted that the inflation method had a higher divergence rate which was 63% even under the correct model ($F=8$) and much higher (86%) for misspecified models ($F=10$). In contrast, the proposed high dimensional EnKF methods had low divergence rates, showing their robustness to the sparsity of the observations.
@@ 6.2 Shallow Water Equation model
The second experiment was conducted on the Shallow Water Equation model (SWE). The SWE is a system of hyperbolic/parabolic partial differential equations governing fluid flow, which is widely used in river flow simulations, as well as in other settings where the water depth is significantly lower than the horizontal length scale of motion. The barotropic nonlinear SWE takes the following form (Lei, Stauffer and Deng, 2012), for $0\leq z_{1}\leq L,0\leq z_{2}\leq D,$
@@
where the fluid velocity components $u$ in $z_{1}$ direction and $v$ in $z_{2}$ direction as well as the fluid depth $h$ are the model variables, $g$ is the gravity acceleration $9.8\mathrm{~{}m}\mathrm{~{}s}^{-2}$, $f$ is the Coriolis parameter defined as constant $10^{-4}\mathrm{~{}s}^{-1}$ using the $f$-plane approximation, $k$ is the diffusion coefficient which was specified as $5\times 10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$, and $L$ and $D$ are the domain limits which were set to $500\mathrm{~{}km}$ and $300\mathrm{~{}km}$, respectively. The SWE model was discretized with a uniform grid spacing of $10\mathrm{~{}km}$ in the $z_{1}$ and $z_{2}$ directions and was solved using the Lax-Wendroff method with a time step of $30$ seconds. Thus, the SWE model had a dimension of $p^{\prime}=50\times 31$ for each state variable $u,v,h$, leading to the state vector $\mathbf{x}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ having dimension $p=3p^{\prime}=4650$. The initial height (depth) field was given by $h(z_{1},z_{2})=H_{0}+H_{1}\tanh\big{(}{9(D/2-z_{2})}/2/D\big{)}+H_{2}\operatorname{sech}^{2}\big{(}9(D/2-z_{2})/{D}\big{)}\sin\big{(}{2\pi}z_{1}/{L}\big{)},$ where $H_{0},H_{1}$ and $H_{2}$ were set to be $50\mathrm{~{}m},5.5\mathrm{~{}m}$ and $3.325\mathrm{~{}m}$, respectively. The initial velocity field was derived from the initial height field with the geostrophic relation.
The SWE was simulated for $48$ hours with a total of 5760 time steps at 30 second intervals after a half hour (60-steps) pre-running. The observations of $u,v$ and $h$ were made at $q^{\prime}=310$ grid points from $10$ randomly selected rows in $z_{1}$ direction every 3 hours, namely 360 time points. The whole observation vector was $\mathbf{y}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ with dimension $q=3q^{\prime}=930$, which means the observation data were quite sparse with $q=0.2p$. To generate the imperfect models, the diffusion coefficient $k$ was also considered as $10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$ in the assimilation process. The observation error variances were specified as $\sigma_{\mathrm{u}}^{2}=\sigma_{\mathrm{v}}^{2}=0.5,\sigma_{\mathrm{h}}^{2}=1.0$ for the three dimensions and the observation error covariance matrix was $\mathbf{R}_{t}=diag(\sigma_{\mathrm{u}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{v}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{h}}^{2}\mathbf{I}_{p^{\prime}})$. The ensemble size $n$ was 100, and the initial ensemble was generated by adding random noises from $\mathrm{N}_{p}(0,\mathbf{I}_{p})$ to the initial state $\mathbf{x}_{1}$.
@@
Figure 3 reports the RMSEs at each assimilation time step of the five EnKF methods ( the standard EnKF, the inflation and the three high dimensional EnKF) to the oracle analysis states $\mathbf{x}_{t}^{a}$ and the true states $\mathbf{x}_{t}$ under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models. From Figure 3, the RMSEs at each assimilation time step of the three proposed high dimensional EnKF schemes became lower and lower along the time for both the correct and misspecified models. While the RMSEs of the standard EnKF increased over time and the inflation method also saw a slight increase in the RMSE for the true model $k=5\times 10^{4}$. Figure 3 shows that the three proposed high dimensional EnKF methods outperformed the inflation method for both the correct and imperfect models with much smaller RMSEs after the system settled down. The three proposed EnKF schemes saw much larger decreases in assimilation RMSEs in comparison to the standard EnKF and the inflation methods and the differences became larger over time, especially under the misspecified model. Among the three high dimensional EnKF schemes, the one utilizing the banding estimators again showed superior performance as in the simulations of the L96 model.
Table 3 summarizes the average analysis RMSEs to the oracle analysis state and true state over the last 24 hours, together with their $25\%$ and $75\%$ quantiles, by the standard EnKF, the inflation and the three high dimensional EnKF methods for both the correct and misspecified SWE models with $p=4650,q=930$. The results of Table 3 reveal that the high dimensional EnKF methods had significantly lower RMSEs and narrower 25%-75% quantile bands when compared to the standard EnKF and the inflation-based EnKF. These results were particularly evident in cases where the model was imperfect. Specifically, the proposed high dimensional EnKF schemes were able to reduce the RMSE to the oracle analysis states by 41% in comparison with the inflation method. This provides evidence to support the notion that the high dimensional EnKF methods could attain more accurate and robust assimilated outcomes for SWE models even when the observations were extremely sparse and the state variables were ultra-high dimensional. This also suggests that the proposed high dimensional EnKF methods are highly appealing, especially considering that geophysical models in earth science often possess an ultra-high dimension.
@@ 7 Sea Temperature Assimilation in Northwest Pacific
Ocean is a complex nonlinear system containing motions in various spatial and temporal scales. The multiscale dynamic processes of the ocean and its cross-scale and cross-sphere interactions determine the material and energy cycles of the ocean, which profoundly affect the ocean’s heat absorption and carbon storage capacity, and regulate the evolution of global climate and marine ecosystems.
@@
In this section, we demonstrate how the proposed method is applied to sea temperature assimilation based on the oceanography model named Regional Oceanic Modeling System (ROMS) with satellite data. Satellites provide a valuable tool for environmental and ocean monitoring. While the satellites produce high-resolution sea surface temperature data, they can not provide ocean measurement beneath the ocean surface where the physical model ROMS plays a key role. ROMS is a three-dimensional, free-surface, terrain-following ocean general circulation model that solves the Reynolds-averaged Navier–Stokes equations using the hydrostatic and Boussinesq assumptions (Haidvogel et al., 2000; Shchepetkin and McWilliams, 2005; Haidvogel et al., 2008). Terrain-following vertical coordinates are employed which allow for greater vertical resolution in regions with complex bathymetry.
@@
The data assimilation was conducted from January 1st, 2008 to March 31st, 2008 over a study area 25.5N-34.7N, 144.4E-154E. The spatial resolution of the ROMS was $0.56^{\circ}\times 0.49^{\circ}$ with $18\times 20$ horizontal grids and $18$ uneven vertical layers over the water column. These made the dimension $p$ for the state variable of sea temperature ($\mathbf{x}_{t}$) to be $6480=18\times 20\times 18$. To scale down the computation, we did not include the other ocean variables namely salinity, horizontal velocity and vertical velocity of ocean current as the assimilated variables, and let them be internally determined by ROMS at each time step with the sea temperature as the assimilated variables.
In our study, the integration step (finer time discretization step) for ROMS was 5 minutes and the assimilation time length was daily which was also the frequency of the observations. The observational data were the daily Optimum Interpolation Sea Surface Temperature (OISST) from satellites provided by National Oceanic and Atmospheric Administration (NOAA), which can be downloaded at https://www.ncei.noaa.gov/products/optimum-interpolation-sst. OISST is a long-term data records for Sea Surface Temperature (SST) that incorporates observations from satellites, ships and buoys into regular global grids, where ships and buoys observations are used to correct satellite SST biases. Descriptions for OISST data can be found in (Reynolds et al., 2007; Banzon and Reynolds, 2013). As the OISST only provides sea surface temperature, the observation dimension was $q=360$.
@@
To run ROMS, we need to input the initial state of sea temperature $\mathbf{x}_{1}$, the boundary conditions and smoothed bathymetry data (sea bottom topography). The bathymetry data came from ETOPO2, which is a global map of the earth’s seafloor and land elevations based on satellite altimetry, shipboard echo-sounding, and other data sources and can be downloaded from https://www.ncei.noaa.gov/products/etopo-global-relief-model, as shown in Figure S.1 of the SM. The initial state ($\mathbf{x}_{1}$) was from the Global Ocean Reanalysis and Simulations (GLORYS) (https://data.marine.copernicus.eu/products) on January 1st. The boundary conditions of ROMS can be found in Marchesiello, McWilliams and Shchepetkin (2001).
@@
As the ocean temperature, salinity and current fields have different spatial scales and resolution in the three dimension study region both in the ROMS and the physical ocean, we transform the study region to a scale down space to facilitate the bandwidth selection in the banding estimator. Specifically, the grid points in the study region are denoted by coordinates $\{(u,v,h),u=1,2,\cdots 18,v=1,2,\cdots 20,h=1,2,\cdots 18\}$ in the transformed space of unit rectangles. The banding estimator is performed in the standardized grids. Specifically, the temperatures on the physical grids are labelled as $\{x_{u,v,h},u=1,2,\cdots 18,v=1,2,\cdots 20,h=1,2,\cdots 18\}$ in the standardized grids and the whole state vector in the physical grids can be labelled by $\mathbf{x}=(x_{1,1,1},\cdots,x_{18,20,18})^{\mathrm{T}}$ with the subscript $t$ hidden. The banding estimator with bandwidth $k$ means that we keep the covariances between $x_{u,v,h}$ and $x_{u^{\prime},{v^{\prime}},h^{\prime}}$ on the standardized grids if they satisfy $\max\{|u^{\prime}-u|,|v^{\prime}-v|,|h^{\prime}-h|\}\leq k$.
Considering that ocean is comparatively stable over a short period and to reduce the computation cost, we selected the bandwidth $k$ in the banding estimator by minimizing the out-of-sample RMSE over the first month of the study period between assimilated temperatures and temperatures recorded by Argo in different depths for four candidate levels $k\in\{1,2,3,4\}$. It was found that the RMSE increased with $k$ and was the smallest at $k=1$, which led to the use of $k=1$ as the bandwidth. The inflation factor for the inflation method was similarly selected by minimizing the out-of-sample RMSE over the first month which was $1.1$.
@@
We performed the high dimensional EnKF scheme using the banding estimator with an ensemble size of $n=30$ to obtain the assimilated three-dimensional sea temperature for the study period and region. We also compared the proposed method with the standard EnKF and the inflation method. The out-of-sample performances of these methods were assessed on data collected under the ocean Argo project over the study period in the study region, which can be accessed at https://argo.ucsd.edu/data/. The Argo project collects information from inside the ocean using floats that initially drift with the ocean currents and then surface up and down between the surface and a mid-sea to collect measurements. As the observations provided by Argo are in-situ observations at various depths, they are more precise than the remote sensing data by the satellites, and hence, are used as substitutes for the ground truth.
Figure 4 displays both the in-sample RMSEs of the assimilated sea surface temperature against the OISST observations, and the out-of-sample RMSEs of the assimilated sea water temperature against the temperatures recorded by Argos at various depths. It also presents the correlations (Corrs) between the assimilated sea surface temperatures and those from OISST, and the Corrs between the assimilated temperatures with those recorded by Argo floats. Figure 4 shows that the assimilation RMSEs to the OISST data of the high dimensional EnKF kept stably at a much lower level during the whole assimilation period compared with those for the standard EnKF and the inflation methods. At the same time, the correlations between the OISST and the assimilated sea surface temperatures by the high dimensional EnKF were much higher and more stable than the two other methods which fluctuated quite substantially over the study period. On the out-of-sample Argo data, the high dimensional EnKF with the banding method produced assimilated sea temperatures which were closer to Argo data as the assimilation RMSEs were slightly smaller and the correlations to Argo data were slightly higher for most days during the assimilation period. These implies that the proposed high dimensional EnKF algorithm could achieve more accurate and robust assimilation results in practice.
Table 4 summarizes the average assimilation RMSEs and the correlations of the assimilated sea temperatures $\hat{\mathbf{x}}_{t}^{a}$ to the OISST data and the Argo data over the whole assimilation period by different methods together with their standard errors (in the parentheses). Obviously, the proposed high dimensional EnKF with banding estimator outperformed the standard EnKF and the inflation methods with much smaller RMSEs and much higher correlations to both the OISST and the Argo data. The reductions in the RMSEs and the increases in the correlations against both the OISST and the Argo data of the proposed high dimensional EnKF algorithm were significant compared with the standard EnKF and the inflation methods and they were much more apparent against the OISST data, suggesting that the high dimensional EnKF could make better use of the observation information. Besides, the standard errors of the RMSEs and correlations of the high dimensional EnKF were also smaller compared to the standard EnKF and the inflation methods, which was apparent from Figure 4. This implies that high dimensional EnKF provided more robust assimilation results.
@@
Figure 4: Assimilation RMSEs and correlations to the OISST data (left panels) and the Argo data (right panels) from Jan. 1st 2008 to Mar. 31st 2008 using the standard EnKF (blue), the modified EnKF methods: inflation (red) and banding (green), respectively. The dashed line represents the average assimilation RMSEs and correlations to the OISST data and the Argo data over the whole assimilation period.
@@
Figure 5 depicts the spatial distribution of the assimilated sea surface temperature over the study area on January 1st, February 15th and March 31st in 2008. From the picture, we can see the spatial distribution of the assimilated sea surface temperature varied over time and that there was an overall downward trend in the study region during the study period.

### v2

@@ 5.1 Lorenz-96 model
Panels (d), (g)-(f) of Figure 2 demonstrate the influence of the ensemble size on the analysis RMSEs. Although the analysis RMSEs of the inflation and the HD-EnKF methods all became smaller with the increase of the ensemble size $n$, the proposed HD-EnKFs outperformed the inflation method, especially when the ensemble size $n$ was small (panels (d) and (g) with $n=20$). This suggests that the proposed methods could provide more accurate analysis states when the ensemble size was not big enough. Ensemble size is a significant restriction for the large geophysical models as they are extremely costly to run.
@@
Table 1: Average RMSEs to the oracle analysis state and true state of the five EnKF schemes and their divergence rates (in parentheses) for Lorenz-96 model with $p=40$, $q=30$, $n=30$, $\sigma_{0}=0.1$, and $F=8$ for the correct and other $F$ values for misspecified models.
@@
Table 1 displays the analysis RMSEs of different methods and their divergence rates under sparser observations $q=30$ while $p=40$, which illustrates the influence of the sparsity in the observations on the assimilation accuracy. Although the analysis RMSEs of different assimilation methods were slightly higher than those under more observations $p=q=40$, the RMSEs of the HD-EnKF methods were smaller than those of the other methods for imperfect models ($F\neq 8$) and were comparable to the inflation method for the correct model ($F=8$). However, it is noted that the inflation method had a higher divergence rate which reached 63% even under the correct model ($F=8$) while the proposed methods had low divergence rates, showing their robustness to the sparsity of the observations.
@@ 5.2 Shallow Water Equation model
The second experiment was conducted on the Shallow Water Equation model (SWE). The SWE is a system of hyperbolic/parabolic partial differential equations governing fluid flow, which is widely used in river flow simulations, as well as in other settings where the water depth is significantly lower than the horizontal length scale of motion. The barotropic nonlinear SWE takes the following form (Lei et al., 2012), for $0\leq z_{1}\leq L,0\leq z_{2}\leq D,$
@@
where the fluid velocity components $u$ in $z_{1}$ direction and $v$ in $z_{2}$ direction as well as the fluid depth $h$ are the model variables, $g$ is the gravity acceleration $9.8\mathrm{~{}m}\mathrm{~{}s}^{-2}$, $f$ is the Coriolis parameter defined as constant $10^{-4}\mathrm{~{}s}^{-1}$, $k$ is the diffusion coefficient specified as $5\times 10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$, and $L$ and $D$ are the domain limits set as $500\mathrm{~{}km}$ and $300\mathrm{~{}km}$, respectively. The SWE was discretized with a uniform grid spacing of $10\mathrm{~{}km}$ in the $z_{1}$ and $z_{2}$ directions and solved by the Lax-Wendroff method with a time step of $30$ seconds. Thus, the SWE model had a dimension of $p^{\prime}=50\times 31$ for each state variable $u,v,h$, leading to the state vector $\mathbf{x}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ of dimension $p=3p^{\prime}=4650$. The initial height (depth) field was given by $h(z_{1},z_{2})=H_{0}+H_{1}\tanh\big{(}{9(D/2-z_{2})}/2/D\big{)}+H_{2}\operatorname{sech}^{2}\big{(}9(D/2-z_{2})/{D}\big{)}\sin\big{(}{2\pi}z_{1}/{L}\big{)},$ where $H_{0},H_{1}$ and $H_{2}$ were set to be $50\mathrm{~{}m},5.5\mathrm{~{}m}$ and $3.325\mathrm{~{}m}$, respectively. The initial velocity field was derived from the initial height field with the geostrophic relation.
The SWE was simulated for $48$ hours with a total of 5760 time steps at 30 second intervals after a half hour (60-steps) pre-running. The observations of $u,v$ and $h$ were made at $q^{\prime}=310$ grid points from $10$ randomly selected rows in each direction every 3 hours, making the whole observation vector $\mathbf{y}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ be of dimension $q=3q^{\prime}=930$, which means the observations were quite sparse with $q=0.2p$. To generate the imperfect models, the diffusion coefficient $k$ was also considered as $10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$ in the assimilation process. The observation error variances were $\sigma_{\mathrm{u}}^{2}=\sigma_{\mathrm{v}}^{2}=0.5,\sigma_{\mathrm{h}}^{2}=1.0$ for the three dimensions and $\mathbf{R}_{t}=diag(\sigma_{\mathrm{u}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{v}}^{2}\mathbf{I}_{p^{\prime}},\sigma_{\mathrm{h}}^{2}\mathbf{I}_{p^{\prime}})$. The ensemble size $n$ was 100, and the initial ensemble was generated by adding random noises from $\mathrm{N}_{p}(0,\mathbf{I}_{p})$ to the initial state $\mathbf{x}_{1}$.
@@
Figure 3 reports the RMSEs at each assimilation time step of the five EnKF methods (the standard EnKF, inflation and three HD-EnKF) to the oracle analysis states $\mathbf{x}_{t}^{a}$ and the true states $\mathbf{x}_{t}$ under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models while Table 2 summarizes the average analysis RMSEs over the last 24 hours (2880 steps), together with their $25\%$ and $75\%$ quantiles. Figure 3 shows the three proposed HD-EnKF methods outperformed the inflation and the standard EnKF methods for both the correct and imperfect models with much smaller RMSEs after the system settled down and that the superiority became more apparent over time, especially under the misspecified model. Specifically, the RMSEs at each assimilation time step of the three proposed HD-EnKF schemes became lower and lower over time compared to increased RMSEs for the standard EnKF for all cases and slight increases RMSEs for the inflation method under true models.
It is obvious from Table 2 that the HD-EnKF methods had significantly lower RMSEs and narrower 25%-75% quantile bands with more evident superiority under imperfect models. Specifically, the proposed HD-EnKFs were able to reduce the RMSE by 41% compared to the inflation method, implying that the proposed methods could attain more accurate and robust assimilated outcomes even under extremely sparse observations and ultra-high dimensional state variables. Thus, the proposed methods are highly appealing, especially considering that geophysical models in earth science often possess ultra-high dimensions.
@@ 6 Sea Temperature Assimilation in Northwest Pacific
This section demonstrates how the proposed method is applied to sea temperature assimilation based on the oceanography model named Regional Oceanic Modeling System (ROMS) with satellite data. ROMS is a three-dimensional, free-surface, terrain-following ocean general circulation model that solves the Reynolds-averaged Navier–Stokes equations using the hydrostatic and Boussinesq assumptions (Shchepetkin and McWilliams, 2005; Haidvogel et al., 2008). While the satellites produce high-resolution sea surface temperature data, they can not provide ocean measurement beneath the ocean surface where the physical model ROMS plays a key role.
@@
The data assimilation was conducted from January 1st, 2008 to March 31st, 2008 over a study area 25.5N-34.7N, 144.4E-154E. The spatial resolution of the ROMS was $0.56^{\circ}\times 0.49^{\circ}$ with $18\times 20$ horizontal grids and $18$ uneven vertical layers over the water column, making the dimension $p$ for the state variable of sea temperature ($\mathbf{x}_{t}$) to be $6480=18\times 20\times 18$. To scale down the computation, we did not include the other ocean variables namely salinity and current velocity as the assimilated variables, and let them be internally determined by ROMS at each step with the sea temperature as the assimilated variables. The assimilation time length was daily which was also the frequency of the observations. The satellite data were the daily Optimum Interpolation Sea Surface Temperature (OISST; https://www.ncei.noaa.gov/products). OISST is a long-term data records for Sea Surface Temperature (SST) that incorporate satellite observations into regular global grids using ships and buoys observations correct satellite SST biases (Banzon and Reynolds, 2013). As the OISST only provides sea surface temperature, the observation dimension was $q=360$.
@@
To run ROMS, we need to input the initial state of sea temperature $\mathbf{x}_{1}$, the boundary conditions and smoothed bathymetry data (sea bottom topography). The bathymetry data came from ETOPO2 (https://www.ncei.noaa.gov/products/etopo-global-relief-model), which is a global map of the earth’s seafloor and land elevations as shown in Figure S.1 of the SM. The initial state ($\mathbf{x}_{1}$) was from the Global Ocean Reanalysis and Simulations (GLORYS; https://data.marine.copernicus.eu/products) on January 1st. Thorough descriptions of boundary conditions of ROMS can be found in Marchesiello et al. (2001).
@@
As the ocean temperature, salinity and current fields have different spatial scales and resolution in the three dimension study region both in the ROMS and the physical ocean, we transform the study region to a scale down space to facilitate the bandwidth selection in the banding estimator. Specifically, the grid points in the study region are denoted by coordinates $\{(u,v,h),u=1,\cdots 18,v=1,\cdots 20,h=1,\cdots 18\}$ in the transformed space of unit rectangles, where the banding estimator is performed. The temperatures on the physical grids are labelled as $\{x_{u,v,h}\}$ in the standardized grids and the whole state vector can be labelled by $\mathbf{x}=(x_{1,1,1},\cdots,x_{18,20,18})^{\mathrm{T}}$ with the subscript $t$ hidden. The banding estimator with bandwidth $k$ means that we keep the covariances between $x_{u,v,h}$ and $x_{u^{\prime},{v^{\prime}},h^{\prime}}$ on the standardized grids if they satisfy $\max\{|u^{\prime}-u|,|v^{\prime}-v|,|h^{\prime}-h|\}\leq k$.
@@
We performed the HD-EnKF using the banding estimator with an ensemble size of $n=30$ to obtain the assimilated three-dimensional sea temperature for the study period and region. We also compared the proposed method with the standard EnKF and the inflation method. The out-of-sample performances of these methods were assessed on data collected under the ocean Argo project over the study period and region, accessed at https://argo.ucsd.edu/data/. The Argo project collects information from inside the ocean using floats that initially drift with the ocean currents and then surface up and down between the surface and mid-sea to collect measurements. As the observations provided by Argo are in-situ observations at various depths, they are more precise than the remote sensing data by the satellites, and hence, are used as substitutes for the ground truth for the out-of-sample assessment. The bandwidth $k$ used in the banding estimator was $k=1$, selected by minimizing the out-of-sample RMSE over the first month of the study period between assimilated temperatures and temperatures recorded by Argo at different depths. The inflation factor for the inflation method was similarly selected, which was $1.1$.
@@
Figure 4: Daily assimilation RMSEs and correlations to the OISST(left panels) and the Argo data (right panels) from Jan. 1st 2008 to Mar. 31st 2008 using the standard EnKF (blue), inflation (red) and HD-EnKF with banding (green), respectively. The dashed horizontal line represents the average assimilation RMSEs and correlations over the whole assimilation period.
@@
Figure 4 displays both the in-sample RMSEs and correlations between the assimilated sea surface temperature and the OISST observations, and the out-of-sample RMSEs and correlation between the assimilated sea water temperature and the temperatures recorded by Argos at various depths. From Figure 4, the HD-EnKF provided much lower and more stable assimilation RMSEs and much higher correlations in the entire assimilation period than those for the standard EnKF and the inflation methods. On the out-of-sample performance on the Argo data, the HD-EnKF produced assimilated sea temperatures with smaller RMSEs and higher correlations both on average and for most days during the assimilation period. These imply that the proposed HD-EnKF algorithm could offer more accurate and robust assimilation results in practice. Detailed numerical results of these methods are reported in Table 3, which summarizes the average assimilation RMSEs and correlations of the assimilated $\hat{\mathbf{x}}_{t}^{a}$ to the OISST and the Argo data over the whole assimilation period. The spatial distribution of the assimilated sea surface temperature over the study area on January 1st, February 15th and March 31st in 2008 is depicted in Figure 5, which varied over time and saw an overall downward trend during the study period.